# Defeaturing.bib

### v1

@@ 6 Numerical experiments
In the following we propose some numerical examples in order to validate the proposed estimator. The (weakly) equilibrated flux reconstruction is computed by solving (23) with $p=1$. Some remarks on the stabilized version (see Problem (40)) are reported for the first numerical example. For the elements $K\in\mathcal{G}_{h}^{s}$, the integrals on the active portion $K\cap\Omega_{s}^{\star}$ are computed by defining proper quadrature rules on a subtriangulation of the active portion of the element itself. The integrals on $\gamma_{\bm{a}}^{\star}$ appearing in (23) are computed by quadrature formulas on the intersection between $\mathcal{A}_{h}$ and the feature edges (see for instance [Boffi_nonmatching] for a review on integration of interface terms in non-matching techniques). The same is done also in the computation of the integrals on $\tilde{\gamma}$ required in the evaluation of the defeaturing component of the estimator.
Let us remark that when no feature is included into the geometry, we have $\mathcal{E}_{K}^{2}=\mathcal{E}_{\sigma}^{2}$ for all $K\in\mathcal{A}_{h}^{s}$ and for all $s\in\mathbb{N}$. Indeed, $\mathcal{E}_{\mathrm{g}}=0$ and, since $f\in Q_{h}(\Omega_{0})$, also $\mathcal{E}_{\mathrm{div}}=0$. Hence, when in the following numerical examples we will compare the performances of the proposed adaptive procedure with standard mesh refinement, we will refer to a standard $h$-adaptivity in which the mesh is refined by the Dörfler strategy reported in Section 5.3 according only to the value of $(\mathcal{E}_{\sigma}^{K})^{2}=||\bm{\sigma}_{h}+\nabla u_{h}^{0}||^{2}_{K}$, $K\in\mathcal{T}_{h}^{s}$. We will refer to the adaptive procedure allowing both for mesh and geometric refinement as combined adaptivity. In the tests that follow, convergence trends are shown with respect to the number $N$ of degrees of freedom of $u_{h}^{\star}$. The numerical component of the estimator is approximated as $\mathcal{E}_{\mathrm{num}}=\big{(}\sum_{K\in\mathcal{A}_{h}}\mathcal{E}_{K}^{2}\big{)}^{\frac{1}{2}}$.
@@ 6.1 Test 1: single internal feature
Figure 7: Test1: Convergence of the total estimator and of the overall error against the number of degrees of freedom ($N$). In blue: feature included at the first iteration; in red: feature never included.
The trend of the overall error and of the total error estimator during the adaptive procedure is reported in Figure 7. The red curves refer to the case in which the feature is not included into the geometry and only mesh refinement is performed, whereas the blue curves are obtained allowing both for mesh and geometric refinement. We can observe that, if the feature is not included, the error and the total estimator reach a plateau. On the contrary, if the feature is included into the geometry the error and the total estimator converge at the expected rate, i.e. as $\mathcal{O}(N^{-\frac{1}{2}})$, where $N$ is the total number of degrees of freedom for the variable $u_{h}^{\star}$. In particular, following the procedure devised in Section 5.3, the feature is included at the very first step, due to the predominance of the defeaturing component of the estimator on the local contributions to the numerical component. Hence, excluding the very first iteration, the whole adaptive procedure is steered by the numerical component of the estimator, as in the case in which the feature is not included. However, when the feature is included, the numerical component of the estimator accounts also for the error introduced by the weak imposition of the Neumann condition on the feature boundary. The feature is included at the very first iteration since it is located in a region in which the gradient of the solution is very steep: hence it is expected to have a strong impact on the accuracy of the solution. The step at which the feature is included could be slightly postponed by decreasing the value of $\alpha_{3}$, within the limit in which the defeaturing component of the estimator does not affect the convergence of the total estimator. As discussed in [BCV2022], the optimal value of $\alpha_{3}$ results to be problem dependent and the parameter is usually fine tuned heuristically.
@@
Figure 8: Test1: Convergence of the total estimator and of its components with respect to the number of degrees of freedom ($N$).
@@ 6.2 Test 2: multiple internal and boundary features
The aim of this second numerical experiment is to test the proposed adaptive procedure on a case characterized by the presence of multiple features, both internal and on the boundary. Let again $\Omega_{0}=(0,1)^{2}$ and let us consider a set =$\mathcal{F}=\{F_{i}\}_{i=1}^{37}$ of 37 features, being all regular polygons inscribed in a circle of radius $\epsilon_{i}$, centered in $\bm{x}_{i}^{C}$, having $n_{i}^{e}$ edges and rotated of an angle $\theta_{i}$ with respect to a reference orientation in which one of the vertexes is located in $\bm{x}_{i}^{C}+[0,\epsilon_{i}]$. In particular, denoting by $\mathcal{U}(a,b)$ a standard uniform distribution between $a$ and $b$, we take $\epsilon_{i}\sim\mathcal{U}(1e-3,5e-2)$ and $\theta_{i}\sim\mathcal{U}(0,2\pi)$. For internal features $\bm{x}_{i}^{C}\sim\mathcal{U}(0.1,0.9)^{2}$, while for boundary features $\bm{x}_{i}^{C}\sim\{0,1\}\times\mathcal{U}(0.1,0.9)$. For both cases the number of edges $n_{i}^{e}$ is chosen from a discrete integer uniform distribution between 4 and 16. The actual data used in the experiments are available in the Appendix.
@@
Figure 12 compares the convergence of the error estimator and of its components in two scenarios: the blue curves refer to the case in which the adaptive strategy proposed in Section 5 is applied with $\alpha_{1}=\alpha_{2}=\alpha_{3}=1$, whereas the red curves are obtained without including any feature, i.e. with standard mesh refinement based on the value of $(\mathcal{E}_{\sigma}^{K})^{2}$. As in the previous numerical test, we can observe how, in this second case, the total estimator rapidly reaches a plateau, due to the fact that the defeaturing component of the estimator remains constant. Allowing instead for both mesh and geometric refinement the total estimator tends to converge as $N^{-\frac{1}{2}}$, i.e. as the numerical component. Figure 13 shows the computational mesh after 6 iterations of the adaptive procedure, both for standard mesh adaptivity and for the combined one. At this stage, if no features are included, the total estimator is reduced of about $27\%$ with respect to its initial value while, if we allow also for geometric adaptivity, $\mathcal{E}$ is reduced of about $50\%$, with only 7 features included. As expected, and as shown in Figure 13(b), these 7 features are the biggest among the ones closest to the bottom left corner, which is coherent also with the results in [BCV2022_arxiv]. The adaptive process is stopped as $5\cdot 10^{3}$ degrees of freedom are reached. At this stage, in the combined strategy, all the features are included (all the features are actually included before iteration 26) and the total estimator has been reduced of about $94\%$, against a relative decrease of $47\%$ in the standard mesh adaptivity case. Figure 14 shows the final meshes, both for standard and combined adaptivity.
@@ 6.3 Test 3: multiple internal features, discontinuous coefficient
$\begin{cases}-\nabla\cdot(\kappa\nabla u)=0&\text{in }\Omega\\ u=g_{\mathrm{D}}&\text{on }\Gamma_{\mathrm{D}}\\ \nabla u\cdot\bm{n}=0&\text{on }\Gamma_{\mathrm{N}}\end{cases}$ (64)
@@
$\mu_{1}(t)=\begin{cases}2&\text{if }t\in[-1,-\frac{1}{2})\\ -t+\frac{3}{2}&\text{if }t\in[-\frac{1}{2},0)\\ \frac{3}{2}&\text{if }t\in[0,\frac{1}{2})\\ -t+2&\text{if }t\in[\frac{1}{2},1],\end{cases}\quad\mu_{2}(t)=\begin{cases}-t&\text{if }t\in[-1,-\frac{1}{2})\\ \frac{1}{2}&\text{if }t\in[-\frac{1}{2},0)\\ -t+\frac{1}{2}&\text{if }t\in[0,\frac{1}{2})\\ 0&\text{if }t\in[\frac{1}{2},1],\end{cases}$
@@
Despite having introduced a coefficient $\kappa\neq 1$, we still compute the (weakly) equilibrated flux reconstruction as a discrete approximation of $\bm{\sigma}\in\bm{H}(\text{div},\Omega^{\star})$ satisfying (17). For this reason we have to slightly modify the definition of the estimator, by setting $(\mathcal{E}_{\sigma}^{K})^{2}=||\bm{\sigma}_{h}+\kappa\nabla u_{h}^{\star}||_{K\cap{\Omega^{\star}}}^{2}$, whereas the other components remain unchanged. The (weakly) equilibrated flux reconstruction is still computed by summing the solutions of patch-local problems in the form of (23). The non-unitary coefficient $\kappa$ affects only the right hand side of the system, whose components need to be redefined as
$L_{\bm{a}}(\bm{v_{h}}):=-(\psi_{\bm{a}}(\kappa\nabla u_{h}^{\star}),\bm{v}_{h})_{{\omega}_{\bm{a}}^{\star}}-\frac{1}{h_{\bm{a}}}\langle\psi_{\bm{a}}g,\bm{v}_{h}\cdot\bm{n}\rangle_{\gamma_{\bm{a}}^{\star}}$
$R_{\bm{a}}^{p}(q_{h})=(\psi_{\bm{a}}f-\nabla\psi_{\bm{a}}\cdot(\kappa\nabla u_{h}^{\star}),q_{h})_{{\omega}_{\bm{a}}^{\star}}+p\langle\psi_{\bm{a}}g,q_{h}\rangle_{\gamma_{\bm{a}}^{\star}},$
@@
In the case of $h$-adaptivity we can observe how the total estimator converges very slowly, steered by the defeaturing component, which remains almost constant. The slight decrease of $\mathcal{E}_{\mathrm{def}}$, in particular in the first iterations, is related to the fact that the solution on the features centered in the singularities undergoes quick changes during the first refinement steps. The final mesh is reported in Figure 17(a), while the final numerical solution is shown in Figure 17(b). In the case of the combined adaptivity we can instead observe how the total estimator tends to converge at the expected rate, completely steered by the numerical component, to which it is almost perfectly overlapped. The sudden decrease of the estimator in the first part of the adaptation is related to the fact that all the features in $\mathcal{F}_{\mathrm{sing}}$ are marked for inclusion during the first iterations. After that, in a few steps in which the mesh is refined close to these features, the singularities are finally excluded from the active mesh. The features in $\mathcal{F}_{\mathrm{rand}}$ are instead added a bit later, but all within iteration 22. The final mesh, along with the corresponding numerical solution is reported in Figure 18(a).
@@ 7 Conclusions
An adaptive strategy based on an equilibrated flux a posteriori error estimator was proposed for defeaturing problems, extending the work in [BCGVV] to the case of trimmed geometries. The focus is kept on Poisson’s problem with Neumann boundary conditions on the feature boundary and, in particular, on the negative feature case. Similarly to [BCV2022_arxiv] the adaptive strategy accounts both for standard mesh refinement and geometric refinement, i.e. for feature inclusion. The resulting adaptive strategy is called combined adaptivity.
@@
The use of a reconstructed flux allows us to never evaluate the numerical flux on the boundary of the features, which is a great advantage for finite element discretizations, in which the gradient of the numerical solution is typically discontinuous across element edges. Furthermore, until no features are included into the geometry, it allows us to sharply bound the numerical component of the error, hence reducing the number of parameters to be fine tuned for the application of the adaptive procedure.
The reliability of the estimator was proven in $\mathbb{R}^{d}$, $d=2,3$ and verified on three different numerical examples in $\mathbb{R}^{2}$, also involving multiple features of different sizes, both internal and on the boundary. The generalization to the positive feature case is left to a forthcoming work, as well as the validation of the proposed adaptive strategy on three-dimensional geometries, this second having an impact mainly on computational aspects.
@@ Acknowledgments
Author Denise Grappein acknowledges to be holder of a Postdoctoral fellowship financed by INdAM (Istituto Nazionale di Alta Matematica) and hosted by the Research Unit of Politecnico di Torino.
Appendix A Appendix

### v2

@@ 6 Numerical experiments
In the following we propose some numerical examples in order to validate the proposed estimator. We focus on the $r=1$ case, i.e. we use linear Lagrange finite elements to compute the primal solution $u_{h}^{\star}$, and we reconstruct the (weakly) equilibrated flux in a Raviart–Thomas space of order 1. Some remarks on the stabilized version (see Problem (34)) are reported for the first numerical example. For the elements $K\in\mathcal{G}_{h}^{s}$, the integrals on the active portion $K\cap\Omega_{s}^{\star}$ are computed by defining proper quadrature rules on a subtriangulation of the active portion of the element itself. The integrals on $\gamma_{\bm{a}}^{\star}$ appearing in (25) are computed by quadrature formulas on the intersection between $\mathcal{A}_{h}$ and the feature edges (see for instance [3] for a review on integration of interface terms in non-matching techniques). The same is done also in the computation of the integrals on $\tilde{\gamma}$ required in the evaluation of the defeaturing component of the estimator.
[…]
@@ 6.1 Test 1: single internal feature
Figure 7: Test1: Convergence of the total estimator and of the overall error against the number of degrees of freedom (NN). In blue: feature included at the first iteration; in red: feature never included.
The trend of the overall error and of the total error estimator during the adaptive procedure is reported in Figure 7. The red curves refer to the case in which the feature is not included into the geometry and only mesh refinement is performed, whereas the blue curves are obtained allowing both for mesh and geometric refinement. We can observe that, if the feature is not included, the error and the total estimator reach a plateau. On the contrary, if the feature is included into the geometry the error and the total estimator converge at the expected rate, i.e. as $\mathcal{O}(N^{-\frac{1}{2}})$, where $N$ is the total number of degrees of freedom for the variable $u_{h}^{\star}$. In particular, following the procedure devised in Section 5.3, the feature is included at the very first step, due to the predominance of the defeaturing component of the estimator on the local contributions to the numerical component. Hence, excluding the very first iteration, the whole adaptive procedure is steered by the numerical component of the estimator, as in the case in which the feature is not included. However, when the feature is included, the numerical component of the estimator accounts also for the error introduced by the weak imposition of the Neumann condition on the feature boundary. The feature is included at the very first iteration since it is located in a region in which the gradient of the solution is very steep: hence it is expected to have a strong impact on the accuracy of the solution. The step at which the feature is included could be slightly postponed by decreasing the value of $\alpha_{3}$, within the limit in which the defeaturing component of the estimator does not affect the convergence of the total estimator. As discussed in [8], the optimal value of $\alpha_{3}$ results to be problem dependent and the parameter is usually fine tuned heuristically.
@@
Figure 8: Test1: Convergence of the total estimator and of its components with respect to the number of degrees of freedom (NN).
@@ 6.2 Test 2: multiple internal and boundary features
The aim of this second numerical experiment is to test the proposed adaptive procedure on a case characterized by the presence of multiple features, both internal and on the boundary. Let again $\Omega_{0}=(0,1)^{2}$ and let us consider a set $\mathcal{F}=\{F_{i}\}_{i=1}^{37}$ of 37 features, being all regular polygons inscribed in a circle of radius $\epsilon_{i}$, centered in $\bm{x}_{i}^{C}$, having $n_{i}^{e}$ edges and rotated of an angle $\theta_{i}$ with respect to a reference orientation in which one of the vertexes is located in $\bm{x}_{i}^{C}+[0,\epsilon_{i}]$. In particular, denoting by $\mathcal{U}(a,b)$ a standard uniform distribution between $a$ and $b$, we take $\epsilon_{i}\sim\mathcal{U}(1e-3,5e-2)$ and $\theta_{i}\sim\mathcal{U}(0,2\pi)$. For internal features $\bm{x}_{i}^{C}\sim\mathcal{U}(0.1,0.9)^{2}$, while for boundary features $\bm{x}_{i}^{C}\sim\{0,1\}\times\mathcal{U}(0.1,0.9)$. For both cases the number of edges $n_{i}^{e}$ is chosen from a discrete integer uniform distribution between 4 and 16. The actual data used in the experiments are available in the Appendix.
@@
Figure 12 compares the convergence of the error estimator and of its components in two scenarios: the blue curves refer to the case in which the adaptive strategy proposed in Section 5 is applied with $\alpha_{1}=\alpha_{2}=\alpha_{3}=1$, whereas the red curves are obtained without including any feature, i.e. with standard mesh refinement based on the value of $(\mathcal{E}_{\sigma}^{K})^{2}$. As in the previous numerical test, we can observe how, in this second case, the total estimator rapidly reaches a plateau, due to the fact that the defeaturing component of the estimator remains constant. Allowing instead for both mesh and geometric refinement the total estimator tends to converge as $N^{-\frac{1}{2}}$, i.e. as the numerical component. Figure 13 shows the computational mesh after 6 iterations of the adaptive procedure, both for standard mesh adaptivity and for the combined one. At this stage, if no features are included, the total estimator is reduced of about $27\%$ with respect to its initial value while, if we allow also for geometric adaptivity, $\mathcal{E}$ is reduced of about $50\%$, with only 7 features included. As expected, and as shown in Figure 13(b), these 7 features are the biggest among the ones closest to the bottom left corner, which is coherent also with the results in [10]. The adaptive process is stopped as $5\cdot 10^{3}$ degrees of freedom are reached. At this stage, in the combined strategy, all the features are included (all the features are actually included before iteration 26) and the total estimator has been reduced of about $94\%$, against a relative decrease of $47\%$ in the standard mesh adaptivity case. Figure 14 shows the final meshes, both for standard and combined adaptivity.
@@ 6.3 Test 3: multiple internal features, discontinuous coefficient
$\begin{cases}-\nabla\cdot(\kappa\nabla u)=0&\text{in }\Omega\\ u=g_{\mathrm{D}}&\text{on }\Gamma_{\mathrm{D}}\\ {\kappa}\nabla u\cdot\bm{n}=0&\text{on }\Gamma_{\mathrm{N}}\end{cases}$ (58)
@@
[…]
@@
Having introduced a coefficient $\kappa\neq 1$, we need to slightly modify the patch-local problems which we solve to obtain our (weakly) equilibrated flux reconstruction. For every $\bm{a}\in{\mathscr{N}_{h}}$ we still solve a problem in the form of (25), with
${m_{\bm{a}}(\bm{\rho}_{h},\bm{v}_{h}):=(\kappa^{-1}\bm{\rho}_{h},\bm{v}_{h})_{{\omega}_{\bm{a}}^{\star}}+\frac{1}{h_{\bm{a}}}\langle\kappa^{-1}\bm{\rho}_{h}\cdot\bm{n},\bm{v}_{h}\cdot\bm{n}\rangle_{\gamma_{\bm{a}}^{\star}}}$
${L_{\bm{a}}(\bm{v_{h}}):=-(\psi_{\bm{a}}\nabla u_{h}^{\star},\bm{v}_{h})_{{\omega}_{\bm{a}}^{\star}}-\frac{1}{h_{\bm{a}}}\langle\kappa^{-1}\psi_{\bm{a}}g,\bm{v}_{h}\cdot\bm{n}\rangle_{\gamma_{\bm{a}}^{\star}}}$
$R_{\bm{a}}(q_{h})=(\psi_{\bm{a}}f-\nabla\psi_{\bm{a}}\cdot(\kappa\nabla u_{h}^{\star}),q_{h})_{{\omega}_{\bm{a}}^{\star}}+\langle\psi_{\bm{a}}g,q_{h}\rangle_{\gamma_{\bm{a}}^{\star}},$
whereas the bilinear form $b_{\bm{a}}$ remains unchanged [34]. We also need to change the definition of the numerical component of the estimator (see e.g. [39]), by setting
[…]
The use of this weighted norm derives from the need of bounding the energy norm of the overall error, which is defined as $\kappa^{\frac{1}{2}}\nabla(u-u_{h}^{\star})||_{\Omega}$.
@@
In the case of $h$-adaptivity we can observe how the total estimator converges very slowly, steered by the defeaturing component, which eventually reaches a plateau. The decrease of $\mathcal{E}_{\mathrm{def}}$ in the first iterations is related to the fact that the solution on the features centered in the singularities undergoes quick changes during the first refinement steps. The final mesh is reported in Figure 17(a), while the final numerical solution is shown in Figure 17(b). In the case of the combined adaptivity we can instead observe how the total estimator tends to converge at the expected rate, completely steered by the numerical component, to which it is almost perfectly overlapped. The sudden decrease of the estimator in the first part of the adaptation is related to the fact that all the features in $\mathcal{F}_{\mathrm{sing}}$ are marked for inclusion during the first iterations. After that, in a few steps in which the mesh is refined close to these features, the singularities are finally excluded from the active mesh. The features in $\mathcal{F}_{\mathrm{rand}}$ are instead added a bit later, but all within iteration 27. The final mesh, along with the corresponding numerical solution is reported in Figure 18(a).
@@ 7 Conclusions
An adaptive strategy based on an equilibrated flux a posteriori error estimator was proposed for defeaturing problems, extending the work in [7] to the case of trimmed geometries. The focus is kept on Poisson’s problem with Neumann boundary conditions on the feature boundary and, in particular, on the negative feature case. Similarly to [10] the adaptive strategy accounts both for standard mesh refinement and geometric refinement, i.e. for feature inclusion. The resulting adaptive strategy is called combined adaptivity.
@@
The use of a reconstructed flux allows us to never evaluate the numerical flux on the boundary of the features, which is a great advantage for finite element discretizations, in which the gradient of the numerical solution is typically discontinuous across element faces. Furthermore, it enables a sharp bound of the numerical component of the error on elements that are not cut by included features, with all contributions involving an unknown constant restricted to cut elements and to the defeaturing component.
The reliability of the estimator was proven in $\mathbb{R}^{d}$, $d=2,3$ and verified on three different numerical examples in $\mathbb{R}^{2}$, also involving multiple features of different sizes, both internal and on the boundary. The generalization to the positive feature case is left to a forthcoming work, as well as the validation of the proposed adaptive strategy on three-dimensional geometries and on higher-order polynomial spaces, these two extensions having an impact mainly on computational aspects.
@@ Acknowledgments
Appendix A Appendix 1
Let us consider the asymmetric patch-local problem: find $(\bm{\sigma}_{h}^{\bm{a}},\lambda_{h}^{\bm{a}})\in\bm{M}_{h}^{\bm{a}}\times Q_{h}^{\bm{a}}$ such that
$\begin{cases}m_{\bm{a}}(\bm{\sigma}_{h}^{\bm{a}},\bm{v}_{h})-b_{\bm{a}}(\bm{v}_{h},\lambda_{h}^{\bm{a}})=L_{\bm{a}}(\bm{v}_{h})&\forall\bm{v}_{h}\in\bm{M}_{h}^{\bm{a}}\\ \hskip 62.59596ptb_{\bm{a}}^{0}(\bm{\bm{\sigma}_{h}^{\bm{a}}},q_{h})=R_{\bm{a}}^{0}(q_{h})&\forall q_{h}\in Q_{h}^{\bm{a}},\end{cases}$ (59)
where $m_{\bm{a}}$, $b_{\bm{a}}$ and $L_{\bm{a}}$ are defined in (21), (22) and (23) respectively, whereas $b_{\bm{a}}^{0}:\bm{M}_{h}^{\bm{a}}\times Q_{h}^{\bm{a}}\rightarrow\mathbb{R}$ and and $R_{\bm{a}}^{0}:Q_{h}^{\bm{a}}\rightarrow\mathbb{R}$ are defined as
$b_{\bm{a}}^{0}(\bm{v}_{h},q_{h}):=(q_{h},\nabla\cdot\bm{v}_{h})_{{\omega}_{\bm{a}}^{\star}}$
$R_{\bm{a}}^{0}(q_{h}):=(\psi_{\bm{a}}f-\nabla\psi_{\bm{a}}\cdot\nabla u_{h}^{\star},q_{h})_{{\omega}_{\bm{a}}^{\star}}.$
For any $\bm{a}\in\mathscr{N}_{h}$ let $c_{\bm{a}}=-\overline{q_{h}}^{\omega_{\bm{a}}}$, with $q_{h}\in Q_{h}({\omega_{\bm{a}}})$. Since $q_{h}+c_{\bm{a}}\in Q_{h}^{\bm{a}}$ we rewrite the second equation in (59) as
According to (27) and (28), and remembering that Neumann boundary condition is imposed in strong form on the whole ${\partial\omega_{\bm{a}}^{\psi}}\cap\Gamma_{\mathrm{N}}^{0}$ we can rewrite (60) as
which holds also for $q_{h}\in\mathcal{P}_{r}(K)$, since the polynomials in $Q_{h}({\omega_{\bm{a}}})$ are discontinuous. We hence have that for any $K\in\mathcal{A}_{h}$ and for any $q_{h}\in\mathcal{P}_{r}(K)$
$(\nabla\cdot\bm{\sigma}_{h},q_{h})_{K\cap{\Omega^{\star}}}=\hskip-8.5359pt\sum_{\bm{a}\in\mathscr{N}_{h}(K)}(\nabla\cdot\bm{\sigma}_{h}^{\bm{a}},q_{h})_{K\cap{\Omega^{\star}}}$ $=(f,q_{h})_{K\cap{\Omega^{\star}}}-\hskip-8.5359pt\sum_{\bm{a}\in\mathscr{N}_{h}(K)}(\bm{\sigma}_{h}^{\bm{a}}\cdot\bm{n}+{\psi_{\bm{a}}}g,c_{\bm{a}})_{\gamma_{\bm{a}}^{\star}}$ (62)
Appendix B Appendix 2